# Distributional and mean Li–Yorke chaos for weighted shifts on Fréchet sequence spaces

### 02 Mar 2026 v1

@@ 4 Mean Li–Yorke chaos
    If $p\in[1,\infty)$, then $B_{w}$ is mean Li–Yorke chaotic if and only if the condition (A) of Theorem 31 and the following condition hold:
    • (B1)
      There exist $m\in\mathbb{N}$ and an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers such that for each $k\in\mathbb{N}$ there are $r:=r(k)\in\mathbb{N}$ and scalars $b_{-r,k},\cdots,b_{r,k}\in\mathbb{K}$ with $\sum_{j=-r}^{r}\left|a_{j,p(k)}b_{j,k}\right|^{p}\neq 0$ and
      $\frac{1}{N_{k}\left(\sum_{j=-r}^{r}\left|a_{j,p(k)}b_{j,k}\right|^{p}\right)^{\frac{1}{p}}}\sum_{i=1}^{N_{k}}\left(\sum_{j=-r}^{r}\left|a_{j-i,m}b_{j,k}w_{j-i}\cdots w_{j-1}\right|^{p}\right)^{\frac{1}{p}}\geq k,$
      where $p(k):=m$, if $1\leq k\leq m$, and $p(k):=k$, if $k>m$.
Examples of shifts that are hypercyclic but not mean Li–Yorke chaotic are already known; see, for instance, [7, Example 23]. In what follows, to illustrate the use of the necessary direction in Theorem 31 (in particular, Corollary 34), we will give an example of a weighted backward shift on $s(\mathbb{Z})$ that is hypercyclic, but it is not mean Li–Yorke chaotic.
@@
$\left\|\frac{e_{j+n_{k}}}{w_{j}\cdots w_{j+n_{k}-1}}\right\|_{\ell}\leq C^{\prime}_{\ell}\frac{\left(\left|j-n_{k}\right|+1\right)^{\ell}}{2^{n_{k}}}\quad\text{for all }k\in\mathbb{N}.$
@@
$\nu=(\nu_{n})_{n\in\mathbb{Z}}:=(\underset{j\leq 0}{\underbrace{\cdots,1,1,1,1}},1,B_{1},1,1,2,B_{2},2,1,\cdots,1,2,\cdots,n,B_{n},n,\cdots,2,1,\cdots),$
@@
$\lim_{N\to\infty}\frac{1}{N}\sum_{k=1}^{N}d(w_{-k}\cdots w_{-1}e_{-k},0)=\lim_{N\to\infty}\frac{1}{N}\sum_{k=1}^{N}\frac{1}{2^{k}}=0,$
then the condition (A) of Theorem 31 is satisfied. Now, for each $k\in\mathbb{N}$ take $N_{k}:=2\sum_{\ell=1}^{k}\ell+\sum_{\ell=1}^{k}10^{\ell}$. Then,
$\frac{1}{N_{k}\left|\nu_{N_{k}}\right|}\sum_{i=1}^{N_{k}}\left|\nu_{N_{k}-i}\right|$ $=\frac{\left(2\sum_{j=1}^{k}\sum_{i=1}^{j}i\right)+\left(\sum_{j=1}^{k}\sum_{i=1}^{10^{j}}(j+1)\right)}{2\sum_{\ell=1}^{k}\ell+\sum_{\ell=1}^{k}10^{\ell}}$
@@
Thus, the conditions (A1) and (B1) of Corollary 34 are satisfied. Therefore, $B_{w}$ is mean Li-Yorke chaotic.
In the following corollary, we show that Corollary 34 generalizes the characterizations of mean Li–Yorke chaos for weighted backward shifts on the spaces $\ell^{p}(\mathbb{Z})$ and $c_{0}(\mathbb{Z})$ obtained in [9]. To this end, recall that an operator $T$ on a Banach space $Y$ is said to be absolutely Cesàro bounded if there exists a constant $C\in(0,\infty)$ such that
@@
This property is closely related to mean Li–Yorke chaos in the context of Banach spaces. In fact, if an operator $T$ is mean Li–Yorke chaotic, then it is not absolutely Cesàro bounded; see, for instance, [4, 8] for further details.
@@
A weighted backward shift $B_{w}$ on $X:=\ell^{p}(\mathbb{Z})$ ($p\in[1,\infty)$) or $X:=c_{0}(\mathbb{Z})$ with nonzero weights is mean Li–Yorke chaotic if and only if it is not absolutely Cesàro bounded and
$\liminf_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}|w_{-n}\cdots w_{-1}|=0.$ (23)
Proof.
It is straightforward to see that condition of Theorem 31 is equivalent to (23). Moreover, the conditions (A1) and (B1) of Corollary 34 are equivalent to the existence of an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers and a sequence $(y_{k})_{k\in\mathbb{N}}\subset c_{00}(\mathbb{Z})$ such that
$\frac{1}{N_{k}}\sum_{j=1}^{N_{k}}\left\|(B_{w})^{n}(y_{k})\right\|\geq k\left\|y_{k}\right\|,\quad\text{for each }k\in\mathbb{N}.$
This is equivalent to saying that $B_{w}$ is not absolutely Cesàro bounded. ∎
@@
Using Theorem 33, one can obtain unilateral versions of Corollaries 34 and 37 by arguments analogous to the corresponding proofs. We leave the details to the reader.

### 06 Apr 2026 v2

@@ 4 Mean Li–Yorke chaos
    there exist $m\in\mathbb{N}$ and an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers such that, for each $k\in\mathbb{N}$, there are $r:=r(k)\in\mathbb{N}$ and scalars $b_{-r,k},\cdots,b_{r,k}\in\mathbb{K}$ with $\left\|\sum_{j=-r}^{r}b_{j,k}e_{j}\right\|>0$ and
    $\frac{1}{N_{k}\left\|\sum_{j=-r}^{r}b_{j,k}e_{j}\right\|}\sum_{i=1}^{N_{k}}\left\|\sum_{j=-r}^{r}b_{j,k}w_{j-i}\cdots w_{j-1}e_{j-i}\right\|\geq k.$
Proof.
$(\Rightarrow)$ By the Lemma 38, $B_{w}$ admits an absolutely mean irregular vector $x:=(x_{n})_{n\in\mathbb{Z}}\in X$. In particular, $\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}\left\|(B_{w})^{k}(x)\right\|=0.$ Take $j_{0}\in\mathbb{Z}$ such that $x_{j_{0}}\neq 0$. Then, by condition (C), we have
$\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}\|x_{j_{0}}w_{j_{0}-k}\cdots w_{j_{0}-1}e_{j_{0}-k}\|\leq\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}\left\|(B_{w})^{k}(x)\right\|=0.$
By an analogous argument to that of Lemma 29 (with the norm replacing the metric), we obtain item (A). Since $\limsup_{N\to\infty}\frac{1}{N}\sum_{j=1}^{N}\|T^{j}x\|=\infty$, then there exists an increasing sequence $(N_{k})_{k\in\mathbb{N}}$ of positive integers such that for each $k\in\mathbb{N}$
$\frac{1}{N_{k}\left\|x\right\|}\sum_{i=1}^{N_{k}}\left\|(B_{w})^{i}(x)\right\|>k+\frac{1}{2}.$
Therefore, using analogous arguments to those used in the $(\Rightarrow)$ part of the proof of Theorem 30—namely, the density of $c_{00}(\mathbb{Z})$ (since $(e_{n})_{n\in\mathbb{Z}}$ is a basis) and the continuity of $B_{w}$—we conclude item (B).
$(\Leftarrow)$ By item (A) and by an argument analogous to that of Lemma 29 (with the norm replacing the metric), we have $\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}\left\|(B_{w})^{k}(e_{i})\right\|=0$, for all $i\in\mathbb{Z}$. Then, taking $X_{0}:=\left\{e_{n}:n\in\mathbb{Z}\right\}$, we obtain item (i) of MLYCC. Moreover, taking $y_{k}:=\sum_{j=-r}^{r}b_{j,k}e_{j}$ for each $k\in\mathbb{N}$, by item (B), we conclude item (ii) of MLYCC. Therefore, by Lemma 38, $B_{w}$ is mean Li–Yorke chaotic. ∎
Examples of shifts that are hypercyclic but not mean Li–Yorke chaotic are already known; see, for instance, [8, Example 23]. In what follows, to illustrate the use of the necessary direction in Theorem 30, we will give an example of a weighted backward shift on $s(\mathbb{Z})$ that is hypercyclic, but it is not mean Li–Yorke chaotic.
@@
$\left\|\frac{e_{j+n_{k}}}{w_{j}\cdots w_{j+n_{k}-1}}\right\|_{\ell}\leq C^{\prime}_{j}\frac{\left(\left|j+n_{k}\right|+1\right)^{\ell}}{2^{n_{k}}}\quad\text{for all }k\in\mathbb{N}.$
@@
$\nu=(v_{n})_{n\in\mathbb{Z}}:=(\underset{j\leq 0}{\underbrace{\cdots,1,1,1,1}},1,B_{1},1,1,2,B_{2},2,1,\cdots,1,2,\cdots,n,B_{n},n,\cdots,2,1,\cdots),$
@@
$\lim_{N\to\infty}\frac{1}{N}\sum_{k=1}^{N}\left\|w_{-k}\cdots w_{-1}e_{-k}\right\|=\lim_{N\to\infty}\frac{1}{N}\sum_{k=1}^{N}\frac{1}{2^{k}}=0,$ (19)
then condition (A) of Theorem 39 is satisfied. Now, for each $k\in\mathbb{N}$ take $N_{k}:=2\sum_{\ell=1}^{k}\ell+\sum_{\ell=1}^{k}10^{\ell}$. Note that $v_{N_{k}}=1$ for all $k\in\mathbb{N}$. Then
$\frac{1}{N_{k}\left\|e_{N_{k}}\right\|}\sum_{i=1}^{N_{k}}\left\|(B_{w})^{i}(e_{N_{k}})\right\|=\frac{1}{N_{k}\left|\nu_{N_{k}}\right|}\sum_{i=1}^{N_{k}}\left|\nu_{N_{k}-i}\right|$ $=\frac{\left(2\sum_{j=1}^{k}\sum_{i=1}^{j}i\right)+\left(\sum_{j=1}^{k}\sum_{i=1}^{10^{j}}(j+1)\right)}{2\sum_{\ell=1}^{k}\ell+\sum_{\ell=1}^{k}10^{\ell}}$
@@
Thus, condition (B) of Theorem 39 is satisfied. Therefore, $B_{w}$ is mean Li-Yorke chaotic.
Remark 42.
If the uncountable set $S$ in Definition 21 can be chosen to be dense, then we say that $f$ is densely mean Li–Yorke chaotic. Moreover, we say that a continuous linear operator $T$ on a Banach space $X$ satisfies the Dense Mean Li–Yorke Chaos Criterion (DMLYCC) if it satisfies conditions (i) and (ii) in item (c) of Lemma 38, with $X_{0}$ being a dense set. Suppose that $X$ is separable, then, by [7, Theorem 21], we have that $T$ is densely mean Li–Yorke chaotic if and only if $T$ satisfies the DMLYCC.
Since in the last example we proved that
$\lim_{N\to\infty}\frac{1}{N}\sum_{k=1}^{N}\left\|w_{-k}\cdots w_{-1}e_{-k}\right\|=0,$
by an analogous lemma to Lemma 29 (with the norm in place of the metric, and the limit instead of the limit inferior), we may conclude that
$\lim_{N\to\infty}\frac{1}{N}\sum_{k=1}^{N}\left\|(B_{w})^{k}(x)\right\|=0,\quad\text{for every }x\in c_{00}(\mathbb{Z}).$
Therefore, we may take $X_{0}$ to be a dense set, and hence $B_{w}$ is in fact densely mean Li–Yorke chaotic.
Recall that an operator $T$ on a Banach space $Y$ is said to be absolutely Cesàro bounded if there exists a constant $C\in(0,\infty)$ such that
@@
This property is closely related to mean Li–Yorke chaos in the context of Banach spaces. In fact, if an operator $T$ is mean Li–Yorke chaotic, then it is not absolutely Cesàro bounded; see, for instance, [4, 7] for further details. Observe that item (B) of Theorem 39 is equivalent to saying that $B_{w}$ is not absolutely Cesàro bounded. Therefore, by specializing Theorem 39 to the spaces $\ell^{p}(\mathbb{Z})$ ($p\in[1,\infty)$) or $c_{0}(\mathbb{Z})$, we recover the characterization given in [9]:
@@
[9, Corollaries 73 and 78] A weighted backward shift $B_{w}$ on $X:=\ell^{p}(\mathbb{Z})$ ($p\in[1,\infty)$) or $X:=c_{0}(\mathbb{Z})$ with nonzero weights is mean Li–Yorke chaotic if and only if it is not absolutely Cesàro bounded and
$\liminf_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}|w_{-n}\cdots w_{-1}|=0.$
@@
We note that the unilateral versions of Proposition 37, Theorem 39 and Corollary 43 can be obtained in a straightforward way by using analogous arguments to those in the proofs of the bilateral versions.